OpenStudy (anonymous):

Write a proportion for the following problem. Eric is making a scale drawing of the ground floor of his house. If the scale is 2 inches = 5 feet, what number of inches will he use for a length of 55 feet?

OpenStudy (anonymous):

22 in

OpenStudy (anonymous):

i got the answer i jus dnt no how to write the problem

OpenStudy (anonymous):

2/5 = 55/x 2/55 = x/5 2/5 = x/55

OpenStudy (anonymous):

one of those

OpenStudy (anonymous):

2/5 = x/55 You had the right idea.